### v1

# Advancing the hBN Defects Database through Photophysical Characterization of Bulk hBN

Abstract
Quantum emitters in hexagonal boron nitride (hBN) have gained significant attention due to a wide range of defects that offer high quantum efficiency and single-photon purity at room temperature. Most theoretical studies on hBN defects simulate monolayers, as this is computationally cheaper than calculating bulk structures. However, most experimental studies are carried out on multilayer to bulk hBN, which creates additional possibilities for discrepancies between theory and experiment. In this work, we present an extended database of hBN defects that includes a comprehensive set of bulk hBN defects along with their excited-state photophysical properties. The database features over 120 neutral defects, systematically evaluated across charge states ranging from –2 to +2 (600 defects in total). For each defect, the most stable charge and spin configurations are identified and used to compute the zero-phonon line, photoluminescence spectrum, absorption spectrum, Huang-Rhys (HR) factor, interactive radiative lifetimes, transition dipole moments, and polarization characteristics. Our analysis reveals that the electron–phonon coupling strength is primarily influenced by the presence of vacancies, which tend to induce stronger lattice distortions and broaden phonon sidebands. Additionally, correlation analysis shows that while most properties are independent, the HR factor strongly correlates with the configuration coordinates. All data are publicly available at https://h-bn.info, along with a new application programming interface (API) to facilitate integration with machine learning workflows. This database is therefore designed to bridge the gap between theory and experiment, aid in the reliable identification of quantum emitters, and support the development of machine-learning-driven approaches in quantum materials research.
1 Introduction
Solid‐state quantum emitters have become essential building blocks for optical quantum technologies, owing to their ability to generate on‐demand single photons that serve as flying qubits in applications such as quantum communication [1, 2], quantum memory [3, 4, 5], quantum computing [6, 7], and quantum sensing [8, 9]. A variety of platforms, including quantum dots [10, 11], nitrogen‐vacancy (NV) centers in diamonds [12, 13], silicon carbide (SiC) [14], two‐dimensional transition‐metal dichalcogenides (2D TMDs) [15, 16], and defects in hexagonal boron nitride (hBN) [17, 18], have demonstrated single-photon emission under different operating conditions. While epitaxial quantum dots exhibit excellent photon purity, they typically require cryogenic operation [10, 11]. Conversely, bulk crystal defects can function at room temperature but suffer from limited photon out-coupling efficiency due to their three‐dimensional geometry [19]. Two-dimensional materials thus offer a compelling compromise. That is, their atomically thin form reduces the total internal reflection; hence, enhancing the out-coupling efficiency.
Among 2D materials, TMD emitters have shown promise, but to date, most demonstrations remain confined to temperatures below 15 K [15, 16]. In contrast, color centers in hBN operate reliably at room temperature, exhibit high quantum efficiency [18], robustness against aging [20, 21], stability [22], and benefit from high out-coupling due to the atomically thin host lattice [17]. In hBN, a single photon is emitted when an optically excited electron relaxes through a two‐level defect state localized within the wide band gap (6 eV) of the material. Given the large band gap, hBN has been proven to accommodate a multitude of defect‐induced mid‐gap states, giving rise to a rich spectrum of optical transitions [23]. Experimentally, quantum emitters in hBN have been observed at photon energies ranging near 1 eV [24], 2 eV [25, 26], 3 eV [27], and 4 eV [28]. This wide range of emission energies makes defect identification a central challenge in hBN research. Currently, only the V center has a well-established match between theory and experiment [29]. For other proposed emitters such as CV [30, 31], CC [28], CCC [32], CCCC [33, 34, 35, 36], OV [37], a clear consensus of all photophysical properties between experiment and theory remains elusive.
This challenge is based on the fact that many different defects can produce similar signatures, such as overlapping zero-phonon line (ZPL) energies, comparable excited-state lifetimes, similar EPR spectra, and Raman features [38]. A major bottleneck in resolving this issue is the absence of a comprehensive and systematic database covering the photophysical properties of a large set of hBN defects. Therefore, robust defect identification requires a broad and detailed characterization of a wide range of properties across many possible defects [35]. To mitigate the high computational demand associated with first-principles calculations, several machine learning (ML) approaches have been developed to predict photophysical properties across large defect spaces [39, 40]. However, these models still rely on DFT benchmarks for validation, as the photophysical properties of hBN defects do not follow simple patterns based on chemical type, symmetry, or defect class. The inherent variability and randomness in defect-related optical properties thus necessitate a comprehensive, DFT-validated repository that includes both ground- and excited-state characteristics to support reliable identification and interpretation of experimental results.
To address this gap, we previously introduced a hBN defect database [41]. However, it was limited to only monolayer hBN and focused primarily on ground-state electronic properties. Since most experimental studies are conducted on bulk hBN samples, there is a pressing need to extend the theoretical framework to suitably accommodate experimental conditions. In this work, we therefore expand the database to include defect properties in bulk hBN and incorporate excited-state photophysical properties. In contrast to our earlier version, as summarized in Fig. 1, the newly developed database includes new defects and an additional analysis of charge stability, where defect charge states ranging from 2 to 2 are systematically evaluated. For each defect, the most stable charge and spin configurations are identified and subsequently used for detailed calculations of both ground-state and excited-state properties. All computed data are compiled into a publicly accessible database available at https://h-bn.info. Importantly, the database now provides comprehensive data for both bulk and monolayer hBN, allowing users to select the most relevant system for their research needs. As machine learning models are becoming increasingly popular in materials research, we also provide an application programming interface (API) for automated data retrieval to support broader usability and facilitate integration with ML workflows. As a result, this work serves as a vital resource for guiding experimental findings, supporting defect identification, and facilitating data-driven approaches in quantum materials research.
2 Method
This section describes the methodology used to generate the data presented in this work. While each method has been described already in Ref. [41], we still review the details here for the sake of completeness. Each block in Fig. 1 corresponds to a step in the computational pipeline and is detailed below.
2.1 DFT calculation details
All first-principles calculations in this work were performed using the Vienna Ab initio Simulation Package (VASP) [42, 43], employing the projector augmented-wave (PAW) method for treating core–valence interactions [44, 45]. To accurately capture the electronic structure of wide-bandgap hBN, we employed the screened hybrid functional HSE06, modified with a Hartree–Fock exchange mixing parameter =0.32, which yields a calculated band gap of 6.09 eV for hBN, in excellent agreement with experiment [46]. A single -point was used for Brillouin-zone sampling in all defect calculations, and a plane-wave cutoff energy of 500 eV was applied. Spin polarization was included throughout to ensure spin conservation in electronic transitions.
Experimentally, bulk hBN can adopt several stacking sequences, most notably AA’ (boron atoms in one layer sit directly above nitrogen atoms in the adjacent layer) and AB (each successive layer is translated by one in‐plane bond length) [47]. In this study, we employ the AA’ configuration exclusively, as total-energy calculations based on hybrid functional identify it to be slightly more stable than the AB stacking [48, 49]. The van der Waals interactions were included via the D3 correction of Grimme et al. [50].
2.2 Selection of investigated defects
To systematically explore defects in bulk hBN, we considered native defects, X, X, XV, XV, XX, and complexes where X is an element from groups III - VI indicated in Fig. 1.
All defect structures were generated within a 662 supercell (288 atoms), as illustrated in Fig. 1, to minimize spurious defect–defect interactions. Ionic positions were relaxed until residual forces fell below 10-2 eV/Å and total energy changes were smaller than 10-4 eV. Note that full relaxation of both atomic positions and lattice parameters was also tested and found to cause zero to 20 nm variation in the ZPL, depending on the defect type. Given this minor effect, we chose to relax only the ionic positions. In total, 124 distinct neutral defect configurations were constructed and carried forward for subsequent charge stability analysis.
2.3 Ground-state properties
2.3.1 Ground-state electronic structures
For hBN defects, the most commonly encountered spin configurations include singlet, doublet, and triplet states, each corresponding to a different number of unpaired electrons. In this work, we determine the preferred spin configuration for each defect by comparing the total energies of all relevant spin states. Only the lowest-energy configuration is selected for subsequent calculations.
2.3.2 Defect formation energy
Defect formation energy is a key quantity used to determine the most stable charge state of a defect under given thermodynamic conditions. For each defect in this study, charge states ranging from 2 to 2 were systematically evaluated by computing their formation energies using the following expression
where is the total energy of the supercell containing the defect in charge state ; is the total energy of pristine hBN; The term represents the number of atoms added or removed from the system, with being a chemical potential of atom species ; The Fermi level corresponds to the electronic chemical potential, referenced to the valence band maximum (VBM) of pristine hBN, denoted by . The final term, , is a correction for spurious electrostatic interactions arising from periodic boundary conditions in charged supercells [51]. In this work, we employed the Spinney Python code for charge correction [52].
By comparing the formation energies across charge states, we identify the most stable charge configuration for each defect, which is then used in all subsequent ground- and excited-state property calculations.
2.4 Excited-state properties
2.4.1 Excited-state electronic structures
To simulate excited-state configurations, we employ the so-called SCF method, a constrained DFT approach in which the occupation of electronic states is manually controlled [53]. In this method, one electron is explicitly promoted to an excited state, allowing us to model the electron excitation.
For spin-polarized systems, such as defects with doublet or triplet ground states, two types of spin-conserving excitations may be possible: spin-up to spin-up or spin-down to spin-down. Both transitions are considered in this work where relevant. In contrast, for singlet defects, only one transition channel is evaluated unless otherwise specified. This ensures consistency with spin-selection rules and captures the dominant optical transitions for each defect configuration.
2.4.2 Zero phonon line (ZPL) calculation
In experimental photoluminescence spectra, the sharpest and most prominent peak is typically associated with the ZPL, which corresponds to a purely electronic transition occurring without the involvement of phonons. In this work, the ZPL is calculated as the total energy difference between the fully relaxed excited-state and ground-state geometries as follows:
While the zero-point vibrational energies (ZPE) can slightly shift the ZPL, typically in the range of a few meV to 0.1 eV, depending on the defect [54, 55], they are neglected in this study under the assumption that ZPE contributions approximately cancel between the ground and excited states. To ensure spin conservation, we treat spin-up and spin-down transitions separately. For certain defects, this results in two distinct spin-conserving excitation pathways, each yielding its own ZPL.
2.4.3 Configuration coordinates
Due to the differences in electronic occupation between the ground and excited states, the atomic geometries of a defect can undergo significant structural relaxation upon excitation. This structural change can be quantified by the so-called configuration coordinate (), which is defined as
where indexes the atoms; denotes the Cartesian components ; and is the mass of atom . The terms and correspond to the equilibrium positions of atom in the ground and excited states, respectively; and is the normalized eigenvector of the phonon mode , indicating the direction of atomic displacement.
2.4.4 HR and DW factors
The Huang-Rhys (HR) factor () can be used to quantify the strength of the electron-phonon coupling. This can be computed from the partial HR factor of each phonon mode , defined as
where is the frequency of the phonon mode ; and is the configuration coordinate displacement corresponding to that mode, as defined previously. Then the HR factor is obtained by summing over all phonon modes, and its spectral distribution can be expressed as
where is the Dirac delta function. In addition to , the Debye–Waller (DW) factor is also computed to characterize the fraction of emission occurring in the ZPL, which is given by DW = .
2.4.5 Absorption and photoluminescence spectra
In this work, the photoluminescence spectrum is calculated from the full-phonon simulation [56], which can be obtained from
where is a normalization constant (typically determined by fitting to experimental data), and is the optical spectral function that captures both ZPL and phonon sidebands, which is computed as
where is a broadening parameter, and is the generating function defined as
Here, is the time-dependent spectral function given by
where represents the phonon-resolved Huang–Rhys factor as described in the previous section.
2.4.6 Transition rate and lifetime
A radiative transition rate quantifies how quickly a transition occurs between two defect states with the same spin polarization. It is given by
where is the radiative transition rate; is the elementary charge; is vacuum permittivity; denotes the ZPL energy; is the modulus square of the dipole moment computed by Eq. (12); and is 1.85, which is the refractive index of the host hBN in the visible [19].
Since the refractive index of hBN can vary between samples or experimental conditions, our database provides an interactive feature that allows users to input a custom refractive index. The transition rate and corresponding radiative lifetime are automatically updated based on the user-defined value. The radiative lifetime () is computed as the inverse of the transition rate
2.4.7 Transition dipole moment and polarization
The polarization characteristics of excitation and emission processes can be derived from the transition dipole moments. To compare theoretical predictions with polarization-resolved PL experiments, both excitation and emission dipoles are rotated by 90∘ and mapped modulo 60∘ to obtain the nearest alignment relative to the hexagonal crystal axes. Although the hexagonal symmetry is 120∘, angles beyond 180∘ are equivalent due to rotational symmetry. To ensure consistency with experimental conditions, dipole vectors are projected onto the -plane of the hBN crystal, allowing the computation of the in-plane polarization visibility.
The transition dipole moments are calculated from the electronic wavefunctions obtained after structural relaxation. These wavefunctions are extracted using the Python package PyVaspwfc [57]. However, since the wavefunctions in the ground and excited states originate from different geometries, two distinct dipole types must be considered: excitation and emission dipoles. The excitation dipole corresponds to the transition from the wavefunction of the optimized ground state to that of the optimized excited state. In contrast, the emission dipole corresponds to the transition from the optimized excited-state wavefunction to the optimized ground-state wavefunction. The transition dipole moment can be obtained from
where are the initial and final wavefunctions; are the corresponding eigenvalues; is the electron mass; and is the momentum operator. Because the involved wavefunctions come from geometrically distinct configurations, a modified version of PyVaspwfc [58] is used to correctly evaluate the dipole moments. The resulting dipole vector may have components in all three Cartesian directions and can be expressed as
If = 0, the transition dipole lies entirely in the -plane, indicating a purely in-plane polarization. This behavior is further analyzed using the linear in-plane polarization visibility, which quantifies the extent to which the emission or absorption is confined within the plane of the hBN crystal.
2.5 Database repository
The hBN defect database is stored as a single SQLite file, hbn_defects_database.db, in the GitHub repository QCS-Theory/hBN-database. It contains a primary table, updated_data, with the following schematics as listed in Tab. 1.
Column name | Type and description | Option |
Host | TEXT: host material identifier | - |
Defect | TEXT: defect chemical formula/code | - |
Defect name | TEXT: descriptive name of the defect | - |
Charge state | INTEGER: integer charge state | - |
Spin multiplicity | TEXT: spin configuration of defects | - |
Optical spin transition | TEXT: allowed optical spin transition | - |
Excitation properties: dipole_x (Debye) | REAL: x-component of excitation dipole moment | abs_dipole_x |
Excitation properties: dipole_y (Debye) | REAL: y-component of excitation dipole moment | abs_dipole_y |
Excitation properties: dipole_z (Debye) | REAL: z-component of excitation dipole moment | abs_dipole_z |
Excitation properties: linear In-plane Polarization Visibility | REAL: Visibility | abs_visibility |
Excitation properties: Intensity (Debye) | REAL: Strength of excitation transition dipole moment | abs_tdm |
Excitation properties: Characteristic time (ns) | REAL: Time of excitation transition | abs_lifetime |
Excitation properties: Angle of excitation wrt the crystal axis | REAL: excitation polarization angle | abs_angle |
Emission properties: dipole_x (Debye) | REAL: x-component of emission dipole moment | ems_dipole_x |
Emission properties: dipole_y (Debye) | REAL: y-component of emission dipole moment | ems_dipole_y |
Emission properties: dipole_z (Debye) | REAL: z-component of emission dipole moment | ems_dipole_z |
Emission properties: linear In-plane Polarization Visibility | REAL: Visibility | ems_visibility |
Emission properties: Intensity (Debye) | REAL: Strength of emission transition dipole moment | ems_tdm |
Emission properties: ZPL (eV) | REAL: ZPL energy | ZPL |
Emission properties: ZPL (nm) | REAL: ZPL wavelength | ZPL_nm |
Emission properties: lifetime (ns) | REAL: Radiative lifetime of emission | lifetime |
Emission properties: Angle of emission wrt the crystal axis | REAL: emission polarization angle | ems_angle |
Emission properties: Polarization misalignment (degree) | REAL: dipole misalignment angle | misalignment |
Emission properties: Configuration coordinate (amu1/2/Å) | REAL: value | Q |
Emission properties: HR factor | REAL: value | HR |
Emission properties: DW factor | REAL: DW value | DW |
Emission properties: Ground-state total energy (eV) | REAL: total energy | E_ground |
Emission properties: Excited-state total energy (eV) | REAL: total energy | E_excited |
Ground-state structure | BLOB: atomic structure file (CIF file) | structure_ground |
Excited-state structure | BLOB: atomic structure file (CIF file) | structure_excited |
Ground-state electronic structure | BLOB: raw OUTCAR file (VASP format) for electronic structure | band_ground |
Excited-state electronic structure | BLOB: raw OUTCAR file (VASP format) for electronic structure | band_excited |
PL lineshape | BLOB: raw PL lineshape file with the broadening parameter () in Eq. (7) equal to 1 | PL |
Raman spectrum | BLOB: raw Raman spectrum file | raman |
To provide access, we use the GitHub REST API contents endpoint. Users type
GET https://raw.githubusercontent.com/ QCS-Theory/hBN-database/main/ hbn_defects_structure.db
This returns the SQLite binary, which users can save locally and open with any SQLite-compatible tool. Alternatively, in Python, one can call get_database(...) function to load and (optionally) filter the data. See Sec. 4 for step-by-step procedures.
3 Results and Discussion
In this section, we first analyze correlations among key photophysical parameters. We then investigate the impact of vacancy on PSB broadening. Finally, we present the overall distribution of hBN defects in a histogram.
3.1 Correlation matrix
The correlation matrix among various photophysical parameters is computed using the Spearman rank correlation method, as shown in Fig. 2. Overall, most parameters exhibit low correlation coefficients, indicating that they are largely independent of one another. However, notable correlations are observed between the configuration coordinate and the HR factor.
These dependencies are expected and can be explained analytically. As shown in Eqs. (4) and (5), the HR factor is proportional to the square of the configuration coordinate, making their correlation inherently strong. This relationship validates the internal consistency of the computed dataset.
3.2 Relationship between the number of vacancies and PSB broadening
We now investigate the behavior of the HR factor, which is directly related to PSB broadening and thus reflects the strength of electron–phonon coupling. In Fig. 3, defects are categorized into four groups based on their vacancy content: no vacancy with a single impurity, no vacancy with complexes, one vacancy, and two vacancies.
Fig. 3(a) shows the correlation between the and the HR factor. As expected from theory, the HR factor increases approximately with the square of , confirming the analytical relationship. Notably, defects without vacancies tend to exhibit lower values of both and the HR factor. In contrast, defects with one or two vacancies span a wider range of values, with a significant fraction showing strong electron–phonon coupling (i.e., higher and HR values). This trend is further supported by the histograms shown in Figs. 3(b) and 3(c), where the distributions of the high HR factor and high mostly belong to the defects with vacancies.
PL lineshape comparisons in Fig. 3(d) further reinforce this observation. Defects without vacancies, including isolated point defects and complexes, typically exhibit the HR factors in the range of 0-2 and 0-3.8, respectively. In contrast, defects containing vacancies show significantly broader HR distributions: 0–12 for single-vacancy defects and 3.5–4 for those with two vacancies. As reflected in the PL lineshapes, vacancy-containing defects tend to display broader PSB. These results suggest that having more vacancies tends to induce stronger lattice distortion upon excitation compared to having no vacancy. This enhanced structural relaxation leads to greater geometric variation between the ground and excited states, thereby increasing the and, consequently, the HR factor.
3.3 Distribution of hBN defects
To further analyze trends across different defect classes, Fig. 4(a) presents the distribution of ZPL energies for defects categorized by the number of vacancies. The results suggest that there is no clear correlation between ZPL energy and the number of vacancies: all three classes (no vacancy, one vacancy, two vacancies) span a wide ZPL range. Interestingly, vacancy-free defects with a single impurity appear to populate both the low and high ends of the ZPL spectrum.
At the lowest ZPL bin, the ON defect is found, with a ZPL energy of 0.24 eV. However, when taking defect formation energies into account, this configuration is energetically less favorable compared to its positively charged counterpart (O), suggesting that such a low-energy transition may not be experimentally observable under typical conditions.
On the opposite end, the GaB defect exhibits a ZPL of 5.82 eV, which is the highest in the dataset. This configuration is thermodynamically stable based on its defect formation energy. However, its electronic structure does not feature well-defined two-level mid-gap states. This behavior can be attributed to the fact that Ga and B belong to the same group in the periodic table, resulting in no difference in the number of valence electrons. Consequently, the electronic transition involved in GaB closely resembles a near-band-edge transition, yielding a ZPL value approaching the intrinsic bandgap of hBN.
For defects containing complexes with one or more vacancies, the ZPL energies consistently fall within the 1–4 eV range. This trend can be attributed to the nature of vacancies, which typically introduce localized electronic states within the mid-gap region of hBN. Given the intrinsic bandgap of approximately 6 eV, these vacancy-induced states often lie near the center of the gap. As a result, optical transitions between such defect levels commonly span energy differences of 1–4 eV, leading to ZPL values confined within this range.
Fig. 4(b) presents the distribution of misalignment between excitation and emission polarizations. The results show that there is no correlation between the misalignment and the number of vacancies. For the majority of defects, excitation and emission dipoles are nearly parallel (minimal misalignment). Since the polarization angles are mapped modulo 60∘ to account for the hexagonal crystal symmetry, most values cluster near 0∘. However, one obvious outlier with a misalignment of 51∘ is identified as the OV defect. Further analysis reveals that this large offset arises from a strong difference in polarization visibility: the excitation dipole is primarily in-plane and linearly polarized, while the emission dipole has an out-of-plane component. As such, the polarization misalignment is substantial.
Lastly, turning to the radiative lifetime distribution shown in Fig. 4(c), the result suggests that most hBN defects exhibit lifetimes ranging from approximately 1 ns to 100 s. Furthermore, there is no clear correlation between the lifetime and the number of vacancies. More broadly, no systematic relationship is observed between the number of vacancies and other key defect properties, as further illustrated in Supplementary S1.
4 Programmatic Retrieval of the hBN Defect Database via the GitHub API
In this section, we describe a procedure for accessing the hBN defect database directly from our GitHub repository using the GitHub REST API, apart from accessing our website https://h-bn.info. We provide a Python module, get_hBN_defects_database.py, which automates (i) accessing the raw SQLite database and (ii) extracting a user‐specified subset of entries. This functionality enables seamless retrieval of the complete dataset, thereby facilitating integration with machine learning workflows. The step-by-step procedure is as follows:
1.
Obtain the retrieval script.
Download get_hBN_defects_database.py and place it in your working directory. -
2.
Import the retrieval function.
In your Python script or interactive session, import the functionfrom get_hBN_defects_database import get_database
3.
Fetch and filter the database.
Invoke get_database() with the desired filtering criteriadata = get_database( option=["ZPL"], host=["monolayer", "bulk"], spin_multiplicity=["singlet", "doublet", "triplet"], charge_state=[-2, -1, 0, 1, 2], optical_spin_transition=["up", "down"], value_range=(2.0, 4.0), download_db=False )
The keyword arguments perform the following functions:
- option
Specifies which database columns to return. The complete set of valid keys is listed in Tab. 1. To retrieve all columns, use
option = ["all"]
- host
selects between the monolayer and bulk hBN datasets. By default, both are returned.
- spin_multiplicity
filters defects by their spin multiplicity. If omitted, all multiplicities are included.
- charge_state
filters defects by charge state. Defaults to all if unspecified.
- optical_spin_transition
filters by optical spin transition (i.e. “up” refers up to up or “down” refers to down to down). Both are returned if this argument is not provided.
- value_range
restricts the numeric range of the selected property. When omitted, no range filtering is applied.
- download_db
If set to True, one can download the raw SQLite database file (named hbn_defects_<options>.db) to the working directory.
4.
Inspect the results.
The object data is a pandas.DataFrame containing the filtered subset, ready for downstream analysis or visualization. If download_db=True, the corresponding .db file will also be saved locally.
The complete code examples are available in Supplementary S2.
5 Further Updates to the Database
In addition to the completely new database of both ground and excited state properties in bulk hBN presented in this work, our hBN defects database has also undergone other updates since its initial release. This includes
1.
Version 1.1: Incorporation of configuration coordinate data for 277 monolayer hBN defects.
2.
Version 1.2: Inclusion of 554 total energy values for each structure, covering both ground- and excited-state configurations for monolayer hBN defects.
3.
Version 1.3: Inclusion of Raman spectra for 100 defects in hBN monolayers obtained from Ref. [38]
6 Conclusion
In summary, we have established an extended database of hBN defects that includes a comprehensive set of defects in bulk hBN, along with their excited-state photophysical properties. In addition to our previous work focusing on monolayer hBN and ground-state characteristics, this new database incorporates additional defect types in bulk hBN, charge-state analysis, and excited-state properties. Key properties such as ZPL, PL lineshape, absorption lineshape, HR factor, interactive radiative lifetimes, transition dipole moments, and polarization characteristics have been systematically computed.
From the current dataset, we computed a correlation matrix and found that only the configuration coordinate and HR factor are strongly correlated, consistent with their analytical relationship. Further analysis showed that defects with vacancies tend to exhibit larger lattice distortions, leading to higher configuration coordinates and HR factors, resulting in broader PSBs. This suggests that broadened PSBs in low-temperature PL may serve as a possible indicator of vacancy-containing defects. Moreover, we analyzed the distribution of all parameters and found no additional strong correlations.
Finally, all data are made publicly available through an online platform https://h-bn.info, alongside the GitHub REST API. This enables programmatic retrieval of the complete SQLite database or user-specified subsets of defect properties. The API returns raw or filtered .db files on demand, and can be seamlessly integrated into automated pipelines and ML workflows. This resource is therefore intended to assist researchers in experimental interpretation, accelerate defect identification, and facilitate data-driven discovery in the field of solid-state quantum emitters.
Data availability
All raw data from this work is available from the authors upon reasonable request. The database with browsing interfaces is freely accessible at https://h-bn.info. The SQLite database file can be downloaded via the GitHub REST API.
Notes
The authors declare no competing financial interest. We acknowledge the assistance of OpenAI’s ChatGPT in managing code snippets. All final decisions on integration and adaptation of those suggestions were made by the authors.
Acknowledgements.
This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy- EXC-2111-390814868 (MCQST) and as part of the CRC 1375 NOA project C2. The authors acknowledge support from the Federal Ministry of Research, Technology and Space (BMFTR) under grant number 13N16292 (ATOMIQS). S.S. acknowledges research funding by Mahidol University (Fundamental Fund FF-111/2568: fiscal year 2025 by the National Science Research and Innovation Fund (NSRF)). The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de) and on its Linux-Cluster.References
- Gisin et al. [2002] N. Gisin, G. Ribordy, W. Tittel, and H. Zbinden, Quantum cryptography, Rev. Mod. Phys. 74, 145 (2002).
- Ahmadi et al. [2024] N. Ahmadi, S. Schwertfeger, P. Werner, L. Wiese, J. Lester, E. Da Ros, J. Krause, S. Ritter, M. Abasifard, C. Cholsuk, and et al., QUICK3 - design of a satellite-based quantum light source for quantum communication and extended physical theory tests in space, Adv. Quantum Technol. 7, 2300343 (2024).
- Reiserer et al. [2016] A. Reiserer, N. Kalb, M. S. Blok, K. J. Van Bemmelen, T. H. Taminiau, R. Hanson, D. J. Twitchen, and M. Markham, Robust Quantum-Network Memory Using Decoherence-Protected Subspaces of Nuclear Spins, Phys. Rev. X 6, 021040 (2016).
- Cholsuk et al. [2024a] C. Cholsuk, A. Çakan, S. Suwanna, and T. Vogl, Identifying electronic transitions of defects in hexagonal boron nitride for quantum memories, Adv. Opt. Mater. 12, 2302760 (2024a).
- Nateeboon et al. [2024] T. Nateeboon, C. Cholsuk, T. Vogl, and S. Suwanna, Modeling the performance and bandwidth of single-atom adiabatic quantum memories, APL Quantum 1, 026107 (2024).
- Ladd et al. [2010] T. D. Ladd, F. Jelezko, R. Laflamme, Y. Nakamura, C. Monroe, and J. L. O’Brien, Quantum computers, Nature 464, 45 (2010).
- Pezzagna and Meijer [2021] S. Pezzagna and J. Meijer, Quantum computer based on color centers in diamond, Applied Physics Reviews 8, 011308 (2021).
- Stern et al. [2024] H. L. Stern, C. M. Gilardoni, Q. Gu, S. Eizagirre Barker, O. F. J. Powell, X. Deng, S. A. Fraser, L. Follet, C. Li, A. J. Ramsay, H. H. Tan, I. Aharonovich, and M. Atatüre, A quantum coherent spin in hexagonal boron nitride at ambient conditions, Nat. Mater. 23, 1379 (2024).
- Gottscholl et al. [2021] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, D. Krauße, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Spin defects in hBN as promising temperature, pressure and magnetic field quantum sensors, Nat Commun 12, 4480 (2021).
- Senellart et al. [2017] P. Senellart, G. Solomon, and A. White, High-performance semiconductor quantum-dot single-photon sources, Nat. Nanotechnol. 12, 1026 (2017).
- Mantynen et al. [2019] H. Mantynen, N. Anttu, Z. Sun, and H. Lipsanen, Single-photon sources with quantum dots in iii-v nanowires, Nanophotonics 8, 747 (2019).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. Hollenberg, The nitrogen-vacancy colour centre in diamond, Physics Reports 528, 1 (2013), the nitrogen-vacancy colour centre in diamond.
- Dutt et al. [2007] M. V. G. Dutt, L. Childress, L. Jiang, E. Togan, J. Maze, F. Jelezko, A. S. Zibrov, P. R. Hemmer, and M. D. Lukin, Quantum Register Based on Individual Electronic and Nuclear Spin Qubits in Diamond, Science 316, 1312 (2007).
- Lienhard et al. [2016] B. Lienhard, T. Schröder, S. Mouradian, F. Dolde, T. T. Tran, I. Aharonovich, and D. Englund, Bright and photostable single-photon emitter in silicon carbide, Optica 3, 768 (2016).
- Klein et al. [2019] J. Klein, M. Lorke, M. Florian, F. Sigger, L. Sigl, S. Rey, J. Wierzbowski, J. Cerne, K. Müller, E. Mitterreiter, P. Zimmermann, T. Taniguchi, K. Watanabe, U. Wurstbauer, M. Kaniber, M. Knap, R. Schmidt, J. J. Finley, and A. W. Holleitner, Site-selectively generated photon emitters in monolayer MoS2 via local helium ion irradiation, Nature Communications 10, 2755 (2019).
- Lei et al. [2023] Y. Lei, Q. Lin, S. Xiao, J. Li, and H. Fang, Optically active telecom defects in mote2 fewlayers at room temperature, Nanomaterials 13, 10.3390/nano13091501 (2023).
- Tran et al. [2016] T. T. Tran, K. Bray, M. J. Ford, M. Toth, and I. Aharonovich, Quantum emission from hexagonal boron nitride monolayers, Nat. Nanotechnol. 11, 37 (2016).
- Nikolay et al. [2019] N. Nikolay, N. Mendelson, E. Özelci, B. Sontheimer, F. Böhm, G. Kewes, M. Toth, I. Aharonovich, and O. Benson, Direct measurement of quantum efficiency of single-photon emitters in hexagonal boron nitride, Optica 6, 1084 (2019).
- Vogl et al. [2019] T. Vogl, M. W. Doherty, B. C. Buchler, Y. Lu, and P. K. Lam, Atomic localization of quantum emitters in multilayer hexagonal boron nitride, Nanoscale 11, 14362 (2019).
- Vogl et al. [2018] T. Vogl, G. Campbell, B. C. Buchler, Y. Lu, and P. K. Lam, Fabrication and deterministic transfer of high-quality quantum emitters in hexagonal boron nitride, ACS Photonics 5, 2305 (2018).
- Liu et al. [2020] W. Liu, Y.-T. Wang, Z.-P. Li, S. Yu, Z.-J. Ke, Y. Meng, J.-S. Tang, C.-F. Li, and G.-C. Guo, An ultrastable and robust single-photon emitter in hexagonal boron nitride, Physica E Low Dimens. Syst. Nanostruct. 124, 114251 (2020).
- Kianinia et al. [2017] M. Kianinia, B. Regan, S. A. Tawfik, T. T. Tran, M. J. Ford, I. Aharonovich, and M. Toth, Robust solid-state quantum system operating at 800 k, ACS Photonics 4, 768 (2017).
- Cholsuk et al. [2022] C. Cholsuk, S. Suwanna, and T. Vogl, Tailoring the emission wavelength of color centers in hexagonal boron nitride for quantum applications, Nanomaterials 12, 2427 (2022).
- Camphausen et al. [2020] R. Camphausen, L. Marini, S. A. Tawfik, T. T. Tran, M. J. Ford, and S. Palomba, Observation of near-infrared sub-poissonian photon emission in hexagonal boron nitride at room temperature, APL Photonics 5, 076103 (2020).
- Dietrich et al. [2018] A. Dietrich, M. Bürk, E. S. Steiger, L. Antoniuk, T. T. Tran, M. Nguyen, I. Aharonovich, F. Jelezko, and A. Kubanek, Observation of fourier transform limited lines in hexagonal boron nitride, Phys. Rev. B. 98, 081414 (2018).
- Hazra et al. [2025] M. Hazra, M. Rieger, A. Kumar, M. N. Mishuk, C. Cholsuk, K. Sripathy, V. Villafañe, K. Müller, J. J. Finley, and T. Vogl, Temperature-dependent emission spectroscopy of quantum emitters in hexagonal boron nitride (2025), arXiv:2507.06783 [physics.optics] .
- Gale et al. [2022] A. Gale, C. Li, Y. Chen, K. Watanabe, T. Taniguchi, I. Aharonovich, and M. Toth, Site-Specific Fabrication of Blue Quantum Emitters in Hexagonal Boron Nitride, ACS Photonics 9, 2170 (2022).
- Mackoit-Sinkevičienė et al. [2019] M. Mackoit-Sinkevičienė, M. Maciaszek, C. G. V. de Walle, and A. Alkauskas, Carbon dimer defect as a source of the 4.1 ev luminescence in hexagonal boron nitride, Appl. Phys. Lett. 115, 212101 (2019).
- Gottscholl et al. [2020] A. Gottscholl, M. Kianinia, V. Soltamov, S. Orlinskii, G. Mamin, C. Bradac, C. Kasper, K. Krambrock, A. Sperlich, M. Toth, I. Aharonovich, and V. Dyakonov, Initialization and read-out of intrinsic spin defects in a van der waals crystal at room temperature, Nature Materials 19, 540 (2020).
- Sajid and Thygesen [2020] A. Sajid and K. S. Thygesen, VC defect as source of single photon emission from hexagonal boron nitride, 2D Mater. 7, 031007 (2020).
- Li and Gali [2022a] S. Li and A. Gali, Bistable carbon-vacancy defects in h-BN, Front. Quantum Sci. Technol. 1, 1007756 (2022a).
- Li et al. [2022] K. Li, T. J. Smart, and Y. Ping, Carbon trimer as a 2 ev single-photon emitter candidate in hexagonal boron nitride: A first-principles study, Phys. Rev. Mater. 6, L042201 (2022).
- Kumar et al. [2023] A. Kumar, C. Cholsuk, A. Zand, M. N. Mishuk, T. Matthes, F. Eilenberger, S. Suwanna, and T. Vogl, Localized creation of yellow single photon emitting carbon complexes in hexagonal boron nitride, APL Mater. 11, 071108 (2023).
- Kumar et al. [2024] A. Kumar, Ç. Samaner, C. Cholsuk, T. Matthes, S. Paçal, Y. Oyun, A. Zand, R. J. Chapman, G. Saerens, R. Grange, S. Suwanna, S. Ateş, and T. Vogl, Polarization dynamics of solid-state quantum emitters, ACS Nano 18, 5270 (2024).
- Cholsuk et al. [2023] C. Cholsuk, S. Suwanna, and T. Vogl, Comprehensive scheme for identifying defects in solid-state quantum systems, J. Phys. Chem. Lett. 14, 6564 (2023).
- Maciaszek and Razinkovas [2024] M. Maciaszek and L. Razinkovas, Blue Quantum Emitter in Hexagonal Boron Nitride and a Carbon Chain Tetramer: a First-Principles Study, ACS Appl. Nano Mater. 7, 18979 (2024).
- Li and Gali [2022b] S. Li and A. Gali, Identification of an oxygen defect in hexagonal boron nitride, J. Phys. Chem. Lett. 13, 9544–9551 (2022b).
- Cholsuk et al. [2025] C. Cholsuk, A. Cakan, V. Deckert, S. Suwanna, and T. Vogl, Raman signatures of single point defects in hexagonal boron nitride quantum emitters, arXiv.org, e-Print Arch., Phys. , arxiv:2502.21118 (2025), arXiv:2502.21118 [cond-mat.mtrl-sci] .
- Shenoy et al. [2020] V. B. Shenoy, N. C. Frey, D. Akinwande, and D. Jariwala, Machine learning-enabled design of point defects in 2D materials for quantum and neuromorphic information processing, ACS Nano 14, 13406 (2020).
- Thiemann et al. [2020] F. L. Thiemann, P. Rowe, E. A. Müller, and A. Michaelides, Machine Learning Potential for Hexagonal Boron Nitride Applied to Thermally and Mechanically Induced Rippling, J. Phys. Chem. C 124, 22278 (2020).
- Cholsuk et al. [2024b] C. Cholsuk, A. Zand, A. Çakan, and T. Vogl, The hBN Defects Database: A Theoretical Compilation of Color Centers in Hexagonal Boron Nitride, J. Phys. Chem. C 128, 12716 (2024b).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficiency of ab-initio total energy calculations for metals and semiconductors using a plane-wave basis set, Comput. Mater. Sci. 6, 15 (1996).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, From ultrasoft pseudopotentials to the projector augmented-wave method, Phys. Rev. B 59, 1758 (1999).
- Elias et al. [2019] C. Elias, P. Valvin, T. Pelini, A. Summerfield, C. J. Mellor, T. S. Cheng, L. Eaves, C. T. Foxon, P. H. Beton, S. V. Novikov, B. Gil, and G. Cassabois, Direct band-gap crossover in epitaxial monolayer boron nitride, Nat. Commun. 10, 2639 (2019).
- Gilbert et al. [2019] S. M. Gilbert, T. Pham, M. Dogan, S. Oh, B. Shevitski, G. Schumm, S. Liu, P. Ercius, S. Aloni, M. L. Cohen, and A. Zettl, Alternative stacking sequences in hexagonal boron nitride, 2D Mater. 6, 021006 (2019).
- Wickramaratne et al. [2018] D. Wickramaratne, L. Weston, and C. G. Van De Walle, Monolayer to Bulk Properties of Hexagonal Boron Nitride, J. Phys. Chem. C 122, 25524 (2018).
- Constantinescu et al. [2013] G. Constantinescu, A. Kuc, and T. Heine, Stacking in bulk and bilayer hexagonal boron nitride, Phys. Rev. Lett. 111, 036104 (2013).
- Grimme et al. [2010] S. Grimme, J. Antony, S. Ehrlich, and H. Krieg, A consistent and accurate ab initio parametrization of density functional dispersion correction (DFT-D) for the 94 elements H-Pu, The Journal of Chemical Physics 132, 154104 (2010).
- Freysoldt et al. [2009] C. Freysoldt, J. Neugebauer, and C. G. Van de Walle, Fully ab initio finite-size corrections for charged-defect supercell calculations, Phys. Rev. Lett. 102, 016402 (2009).
- Arrigoni and Madsen [2021] M. Arrigoni and G. K. Madsen, Spinney: Post-processing of first-principles calculations of point defects in semiconductors with Python, Computer Physics Communications 264, 107946 (2021).
- Jones and Gunnarsson [1989] R. O. Jones and O. Gunnarsson, The density functional formalism, its applications and prospects, Rev. Mod. Phys. 61, 689 (1989).
- Benedek et al. [2023] Z. Benedek, R. Babar, Á. Ganyecz, T. Szilvási, Ö. Legeza, G. Barcza, and V. Ivády, Symmetric carbon tetramers forming spin qubits in hexagonal boron nitride, npj Computational Materials 9, 187 (2023).
- Mishra and Bhattacharya [2019] H. Mishra and S. Bhattacharya, Giant exciton-phonon coupling and zero-point renormalization in hexagonal monolayer boron nitride, Physical Review B 99, 165201 (2019).
- Alkauskas et al. [2014] A. Alkauskas, Q. Yan, and C. G. Van de Walle, First-principles theory of nonradiative carrier capture via multiphonon emission, Phys. Rev. B 90, 075202 (2014).
- Liu [1 01] L. Liu, Pyvaspwfc. https://github.com/liming-liu/pyvaspwfc (accessed 2022-11-01).
- Davidsson [2020] J. Davidsson, Theoretical polarization of zero phonon lines in point defects, J. Phys.: Condens. Matter 32, 385502 (2020).

### v2

@@
 In this work, we employed the Spinney Python code for charge correction [52].
+Although charge states greater than 3 or less than 3 may be possible for certain defects, they were not considered in this work.
+Future studies may extend the analysis to include these regimes in order to identify additional stable defects and potential quantum emitters.
 By comparing the formation energies across charge states, we identify the most stable charge configuration for each defect, which is then used in all subsequent ground- and excited-state property calculations.
@@
 Here, is the time-dependent spectral function given by
-where represents the phonon-resolved Huang–Rhys factor as described in the previous section.
+where represents the phonon-resolved Huang–Rhys factor as described in the previous section. The convergence test of the supercell size is shown in Supplementary S1.
 2.4.6 Transition rate and lifetime
@@
 To provide access, we use the GitHub REST API contents endpoint. Users type
-GET https://raw.githubusercontent.com/ QCS-Theory/hBN-database/main/ hbn_defects_structure.db
+GET https://raw.githubusercontent.com/
+QCS-Theory/hBN-database/main/
+hbn_defects_structure.db
 This returns the SQLite binary, which users can save locally and open with any SQLite-compatible tool.
@@
 Furthermore, there is no clear correlation between the lifetime and the number of vacancies.
-More broadly, no systematic relationship is observed between the number of vacancies and other key defect properties, as further illustrated in Supplementary S1.
+More broadly, no systematic relationship is observed between the number of vacancies and other key defect properties, as further illustrated in Supplementary S2.
+Full example calculations of defect properties in the database are given in Supplementary S3.
 4 Programmatic Retrieval of the hBN Defect Database via the GitHub API
@@
 Import the retrieval function.
-In your Python script or interactive session, import the functionfrom get_hBN_defects_database import get_database
+In your Python script or interactive session, import the functionfrom get_hBN_defects_database import get_database -
 3.
 Fetch and filter the database.
-Invoke get_database() with the desired filtering criteriadata = get_database( option=["ZPL"], host=["monolayer", "bulk"], spin_multiplicity=["singlet", "doublet", "triplet"], charge_state=[-2, -1, 0, 1, 2], optical_spin_transition=["up", "down"], value_range=(2.0, 4.0), download_db=False )
-The keyword arguments perform the following functions:
+Invoke get_database() with the desired filtering criteriadata = get_database( option=["ZPL"], host=["monolayer", "bulk"], spin_multiplicity=["singlet", "doublet", "triplet"], charge_state=[-2, -1, 0, 1, 2], optical_spin_transition=["up", "down"], value_range=(2.0, 4.0), download_db=False )The keyword arguments perform the following functions:
 - option
 Specifies which database columns to return. The complete set of valid keys is listed in Tab. 1. To retrieve all columns, use
-option = ["all"]
-- host
+option = ["all"] - host
 selects between the monolayer and bulk hBN datasets. By default, both are returned.
@@
 The object data is a pandas.DataFrame containing the filtered subset, ready for downstream analysis or visualization. If download_db=True, the corresponding .db file will also be saved locally.
-The complete code examples are available in Supplementary S2.
+The complete code examples are available in Supplementary S4.
 5 Further Updates to the Database
@@
 Moreover, we analyzed the distribution of all parameters and found no additional strong correlations.
+We stress that the primary goal of our work is to construct a large database covering comprehensive properties of many different defects.
+This is to serve as a screening tool that helps rule out unlikely defects and narrow down promising candidates.
+This work does not take into account symmetry reduction effects such as the Jahn–Teller distortion; therefore, the specific values of the ZPL may be shifted.
+Such effects should be carefully considered when determining a defect as a promising candidate.
+For instance, to use our database, if an experiment reports a ZPL at 2.2 eV, one should focus on defects with calculated ZPLs between 2.0 and 2.4 eV, taking into account the finite accuracy of DFT calculations.
+This approach effectively eliminates many candidates and suggests the most plausible ones.
 Finally, all data are made publicly available through an online platform https://h-bn.info, alongside the GitHub REST API.
@@
 Holleitner, Site-selectively generated photon emitters in monolayer MoS2 via local helium ion irradiation, Nature Communications 10, 2755 (2019).
-- Lei et al. [2023] Y. Lei, Q. Lin, S. Xiao, J. Li, and H. Fang, Optically active telecom defects in mote2 fewlayers at room temperature, Nanomaterials 13, 10.3390/nano13091501 (2023).
+- Lei et al. [2023] Y. Lei, Q. Lin, S. Xiao, J. Li, and H. Fang, Optically Active Telecom Defects in MoTe2 Fewlayers at Room Temperature, Nanomaterials 13, 1501 (2023).
 - Tran et al. [2016] T. T. Tran, K. Bray, M. J. Ford, M. Toth, and I. Aharonovich, Quantum emission from hexagonal boron nitride monolayers, Nat. Nanotechnol. 11, 37 (2016).
@@
 - Mishra and Bhattacharya [2019] H. Mishra and S. Bhattacharya, Giant exciton-phonon coupling and zero-point renormalization in hexagonal monolayer boron nitride, Physical Review B 99, 165201 (2019).
-- Alkauskas et al. [2014] A. Alkauskas, Q. Yan, and C. G. Van de Walle, First-principles theory of nonradiative carrier capture via multiphonon emission, Phys. Rev. B 90, 075202 (2014).
+- Alkauskas et al.
+[2014] A.
+Alkauskas, B.
+B.
+Buckley, D.
+D.
+Awschalom, and C.
+G.
+Van De Walle, First-principles theory of the luminescence lineshape for the triplet transition in diamond NV centres, New J.
+Phys.
+16, 073026 (2014).
 - Liu [1 01] L. Liu, Pyvaspwfc. https://github.com/liming-liu/pyvaspwfc (accessed 2022-11-01).